OpenStudy (anonymous):

25) One canned juice drink is 15% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 12% orange juice.

OpenStudy (amistre64):

x@15% = 10(12% - 5%)/(15% - 5%)

OpenStudy (anonymous):

let the quantity of 15% juice = x L then qty of 5% juice = (10-x)L 0.15x + 0.05(10-x) = 0.12 * 10 0.15x + 0.5 - 0.05x = 1.2 0.1x = 1.2 - 0.5 x = 0.7/0.1 x = 7 so 7 litres of 15% juice and 3 litres of 5% juice have to be mixed to get 12% juice

OpenStudy (anonymous):

The idea behind all of the questions you are posting, is that they are systems of linear equations. You can also express the problem as follows: 0.15*x + 0.05*y = 0.12*10 x + y = 10

OpenStudy (anonymous):

yes, if they r to be solved by system of linear equations then the method suggested by Alchemista is correct......

OpenStudy (amistre64):

i got a sneaking suspicion that the idea behind all the questions is to get somebody else to give the answers ....

OpenStudy (anonymous):

I don't understand though. I mean 2x2 systems are incredibly easy and all of these problems boil down to exactly that. All you need to do is decipher the English.

OpenStudy (anonymous):

yeah, I have solved exactly the same questions at least 3-4 times already for different people...